# Enhancing Knowledge Distillation of Large Language Models through Efficient Multi-Modal Distribution Alignment

### v1

@@ Appendix C Analysis of Computational Efficiency
We list the elapsed time for distillation training two epochs on GSM8K for some of the distillation targets, as shown in Table 4. Based on the results, we can find that the introduction of ranking loss has a very minimal impact on computational burden, which can be ignored. Additionally, when using only ranking loss, the computational efficiency is improved by eliminating the need for $softmax$ operation on output logits.
@@ Appendix D Analysis of Hyperparameter
And for downstream tasks, we recommend taking the range $k$ to be 10-15 when using ranking loss alone, and $k$ to be 5 for mixing with other distillation targets such as KL. Because when mixing losses, other distillation objectives can bring the two distributions closer together on a broader scale, and a smaller range $k$ helps the ranking loss focus more on peak prediction to achieve better performance. In addition, small changes to this range have a small effect on the results, and dynamic value of the range tends to create a computational burden during batch calculations, so we do not dynamically adjust this hyperparameter. We also present the results on GSM8K in Table 5 after applying different $k$ values to partial distillation objectives, to further validate the above conclusions.
@@
For the losses allocation method, we have found in experiments that the ratio of ranking loss to other distillation targets is better when the ratio is 1 to 3. Besides, small changes have little effect on the KD effect, so we suggest to adopt this lossed allocation directly, as $\mathcal{L}_{\text{total}}=2\cdot\mathcal{L}_{\text{Ranking}}+\mathcal{L}_{\text{logits}}$. As shown in Figure 4, $\mathcal{L}_{\text{Ranking}}$ and $\mathcal{L}_{\text{logits}}$ align two distribution from different perspectives, thus we believe static allocation is appropriate.
@@
Experimental results show that our dynamic allocation strategy achieves better scores on most tasks, further improving the effectiveness of distillation training. Other distillation targets do not show head or tail bias, and in experiments it is found that fixed ratios of losses work better, so there is no need for dynamic losses allocation.
@@ Appendix E Supplementary Experiments in the Pre-Training Task
Results The results in Table 7 show that our method also effectively improves the consistency of peak predictions between the student model and the teacher model on out-of-distribution data during the KD process. This result demonstrates the generalization ability of our method and complements the experimental results in the main text.

### v2

@@ Appendix C Analysis of Computational Efficiency
We list the elapsed time for distillation training two epochs on GSM8K for some of the distillation objectives, as shown in Table 4. Based on the results, we can find that the introduction of ranking loss has a very minimal impact on computational burden, which can be ignored. Additionally, when using only ranking loss, the computational efficiency is improved by eliminating the need for $softmax$ operation on output logits.
@@ Appendix D Analysis of Hyperparameter
And for downstream tasks, we recommend taking the range $k$ to be 10-15 when using ranking loss alone, and $k$ to be 5 for mixing with other distillation objectives such as KL. Because when mixing losses, other distillation objectives can bring the two distributions closer together on a broader scale, and a smaller range $k$ helps the ranking loss focus more on peak prediction to achieve better performance. In addition, small changes to this range have a small effect on the results, and dynamic value of the range tends to create a computational burden during batch calculations, so we do not dynamically adjust this hyperparameter. We also present the results on GSM8K in Table 5 after applying different $k$ values to partial distillation objectives, to further validate the above conclusions.
@@
For the losses allocation method, we have found in experiments that the ratio of ranking loss to other distillation objectives is better when the ratio is 1 to 3. Besides, small changes have little effect on the KD effect, so we suggest to adopt this lossed allocation directly, as $\mathcal{L}_{\text{total}}=2\cdot\mathcal{L}_{\text{Ranking}}+\mathcal{L}_{\text{logits}}$. As shown in Figure 4, $\mathcal{L}_{\text{Ranking}}$ and $\mathcal{L}_{\text{logits}}$ align two distributions from different perspectives, thus we believe static allocation is appropriate.
@@
Experimental results show that our dynamic allocation strategy achieves better scores on most tasks, further improving the effectiveness of distillation training. Other distillation objectives do not show head or tail bias, and in experiments it is found that fixed ratios of losses work better, so there is no need for dynamic losses allocation.
@@ Appendix E Supplementary Experiments in the Pre-Training Task
Results The results in Table 7 show that our method also effectively improves the consistency of peak predictions between the student model and the teacher model on out-of-distribution data during the KD process. These results demonstrate the generalization ability of our method and complement the experimental results in the main text.